OpenStudy (anonymous):

I NEED HELP< I ALMOST GOT IT I THINK>> Which ordered pair is the solution to the system of equations? 2x - y = -5 x + 3y = 22 A. (1, 7) B. (4, 9) C. (7, 19) D. (8, 21)

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Did you get an answer?

OpenStudy (anonymous):

I was leaning towards D, because you just pop in an rdered pair, right?

OpenStudy (igreen):

Yes, you could do that.

OpenStudy (anonymous):

Okay, thanks

OpenStudy (igreen):

But we can re-order the second equation: x + 3y = 22 Subtract 3y to both sides: x = -3y + 22 Now we can plug in -3y + 22 for x in the first equation: 2x - y = -5 2(-3y + 22) - y = -5 Distribute 2 into the parenthesis: -6y + 44 - y = -5 Subtract: -7y + 44 = -5 Subtract 44 to both sides: -7y = -49 Divide -7 to both sides, what's -49 / -7?

OpenStudy (anonymous):

Oh, sorry, negative6 or 7, sorry, I just woke up lol

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Um..-49 / -7 isn't gonna b a negative.

OpenStudy (igreen):

*be

OpenStudy (anonymous):

Oh, yeah.. Its postitive 7 or 6..

OpenStudy (igreen):

Yes, it's positive 7.

OpenStudy (anonymous):

Ok

OpenStudy (igreen):

So 7 is the y-value in the solution. And there's only one option with a y-value of 7..and what's that?

OpenStudy (anonymous):

A.

OpenStudy (igreen):

Yep, so that's our answer.

OpenStudy (anonymous):

Thanks again!